OpenStudy (rainbow_dash):

Find the first three iterates of the given function for the given initial value. f(x) = 2/x, x_0 = 1. I know this seems simple but whenever i click the "simple" answer, it's always wrong :/

OpenStudy (anonymous):

If I rewrite it, is it: x_2 = 2/(x_1)

OpenStudy (rainbow_dash):

how do you figure?

OpenStudy (anonymous):

\[f(x)=\frac{2}{x}\] and \(x_0=1\) right?

OpenStudy (anonymous):

so \(x_1=f(x_0)=f(1)=\frac{2}{1}=2\) for the first one

OpenStudy (anonymous):

next one, replace \(x\) by \(x_1=2\) and get \[x_2=f(x_1)=f(2)=\frac{2}{2}=1\] so \(x_2=1\)

OpenStudy (anonymous):

now that we see we are back were we started from, looks like the patter should be clear, \[x_0=1,x_1=2,x_2=1,x_3=2,x_4=1,...\]

OpenStudy (anonymous):

*pattern (although "patter" sounds good too)